# Can LLM Generate Culturally Relevant Commonsense QA Data? Case Study in Indonesian and Sundanese

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly being used to generate synthetic data for training and evaluating models. However, it is unclear whether they can generate a good quality of question answering (QA) dataset that incorporates knowledge and cultural nuance embedded in a language, especially for low-resource languages. In this study, we investigate the effectiveness of using LLMs in generating culturally relevant commonsense QA datasets for Indonesian and Sundanese languages. To do so, we create datasets for these languages using various methods involving both LLMs and human annotators. Our experiments show that the current best-performing LLM, GPT-4 Turbo, is capable of generating questions with adequate knowledge in Indonesian but not in Sundanese, highlighting the performance discrepancy between medium- and lower-resource languages. We also benchmark various LLMs on our generated datasets and find that they perform better on the LLM-generated datasets compared to those created by humans. [^1] All datasets and codes in this work are available in https://github.com/rifkiaputri/id-csqa. The dataset is available under the Creative Commons Non-Commercial (CC BY-NC 4.0).
Can LLM Generate Culturally Relevant Commonsense QA Data? Case Study in Indonesian and Sundanese
@@ 1 Introduction
The development of Large Language Models (LLMs) is significantly impacting NLP, leading to an increasing trend in the automated generation of datasets, particularly for Question Answering (QA) tasks. However, a major challenge arises with non-English languages, especially those that are underrepresented, like Indonesian and Sundanese, due to the need for cultural context. For the generated data to be fully useful, it must not only be linguistically accurate, but it also needs to reflect the cultural nuances, historical references, and social norms. It is not yet clear whether current LLMs can create QA data that adequately includes the cultural nuances specific to languages such as Indonesian and Sundanese.
@@
In this study, we aim to investigate how well current LLMs generate a commonsense QA dataset that is both linguistically accurate and culturally relevant to Indonesia. We focus on Indonesian, the lingua franca of Indonesia, and Sundanese, one of the local languages in Indonesia. To sum up, our contributions are as follows:
  • We create a new Indonesian and Sundanese CommonsenseQA dataset using various methods, including adapting existing English data Talmor et al. (2019) and generating new datasets manually and automatically as illustrated in Figure 1. The dataset contains approximately 3K LLM-generated and 1.5K human-generated question-answer pairs per language, making it the largest of its kind. For human-generated data, we ensure cultural diversity by involving annotators from various regions, including Java and Bali.
  • We perform a comprehensive analysis to assess the effectiveness of LLMs in creating a culturally relevant commonsense QA dataset. We find that adaptation from English data is less effective, particularly for Sundanese. However, GPT-4 Turbo can generate questions with some basic local knowledge in Indonesian when provided with human-created categories and concepts. Yet, for Sundanese, questions often have major errors.
  • We evaluate several LLMs using our dataset and find that they perform better on the dataset generated by LLMs than humans. This finding reveals the limitation of the synthetic dataset that it is not as challenging for LLMs, especially for larger models such as GPT-4.
@@ 2 Related Work
Datasets and benchmarks for commonsense reasoning are crucial for developing models that understand and reason about real-world complexities. Datasets like COPA Roemmele et al. (2011), X-COPA Ponti et al. (2020), and The Winograd Schema Challenge Levesque et al. (2012) evaluate causal reasoning in real-world scenarios. Other datasets like ARC Clark et al. (2018), OpenBookQA Mihaylov et al. (2018), and Mcscript Ostermann et al. (2018) test commonsense reasoning via question-answering, but some questions require grade-school science knowledge. CommonsenseQA Talmor et al. (2019) presents the type of “purely” commonsense QA in the form of multiple-choice questions based on a concept knowledge from ConceptNet Speer et al. (2017), built through crowdsourcing. One question corresponds to one target concept, and other related concepts become distractors, acting as parts of the answer choices. However, since the dataset is crowdsourced, the crowdworkers bring their cultural background as part of their common knowledge, resulting in many questions that are “commonsense” only in the Western culture.
@@
Various datasets are used to assess language models’ ability to understand cultural nuances specific to a language. FORK Palta and Rudinger (2023) explores culinary cultural biases and assumptions. However, this dataset is relatively small, comprising only 184 questions and focusing solely on US, Indian, and Chinese customs. In the context of Indonesian, IndoMMLU Koto et al. (2023) includes questions from Indonesian exams and covers regional cultural topics, such as the Minangkabau or Sundanese cultures. COPAL-ID Wibowo et al. (2023) is an Indonesian COPA-style dataset written by native speakers, thus incorporating more Indonesian cultural nuances compared to X-COPA. COPAL-ID is composed of approximately 500 questions, with a primary emphasis on the cultural aspects of the Jakarta region. In contrast, our dataset is much bigger and covers a broader range of annotators from various regions of Java and Bali, in addition to Jakarta. We also include Sundanese, addressing the gap in this low-resource language.
@@ 3.1 Commonsense QA: Definition and Scope
The Cambridge Dictionary defines commonsense as "the basic level of practical knowledge and judgment that we all need to help us live in a reasonable and safe way." van Holthoon and Olson (1987) describe commonsense as "knowledge, judgment, and taste which is more or less universal and which is held more or less without reflection or argument." Both definitions highlight commonsense as the practical knowledge in daily life. Talmor et al. (2019) did not explicitly define the scope of commonsense in their QA dataset. Nevertheless, they outlined various skills required to answer the questions in their dataset correctly, such as spatial, cause & effect, social, activity, and many more. To build our dataset, we follow similar skills but focus on adding common or cultural knowledge specific to Indonesian and Sundanese contexts to reduce bias towards Western culture. Generally, one data point is a triple of concept, question, and options (with one correct answer). The data samples are shown in Figure 1.
@@ 3.2 Languages in Indonesia
Indonesia is one of the most culturally and linguistically diverse countries in the world, with more than 700 languages spoken across the country Aji et al. (2022); Eberhard et al. (2021). Of the many languages in Indonesia, Indonesian is a unifying language used nationally. It utilizes the Latin script and was developed from literary "Classical Malay" of the Riau-Johor sultanate Sneddon (2003), with regional variations. Over 80% of Standard Malay’s vocabulary is similar to Indonesian.
Apart from Indonesian, regional languages like Sundanese are spoken by people of the same ethnicity. Sundanese is primarily spoken in West Java, as well as Banten, Jakarta, Central Java, and more. [^2] https://petabahasa.kemdikbud.go.id/infobahasa2.php?idb=56&idp=Jawa%20Barat Regional languages including Sundanese have influenced the formation and development of the Indonesian language. Both languages share similarities, such as their grammatical structure, but also differ significantly in aspects like the number of vowels and morphological features, including affixes. We chose to study the national language and one regional language to illustrate the differences in the commonsense QA data generated via LLMs.
@@ 4 Data Generation Methods
To investigate whether LLMs can generate culturally relevant commonsense QA data in Indonesian and Sundanese, we construct a dataset using various methods with LLMs as data generators. In addition, we employ humans to generate data for comparison. As illustrated in Figure 1, we apply three dataset generation methods: (1) Automatic Data Adaptation, where we leverage LLMs to automatically adapt English CommonsenseQA data to our target languages; (2) Manual Data Generation, where we ask native-speaker human annotators to manually construct the dataset; and (3) Automatic Data Generation, where we utilize LLMs to generate data based on the list of categories and concepts used in method (2).
@@ 4.1 Automatic Data Adaptation
We build the first LLM-generated data by adapting the CommonsenseQA dataset Talmor et al. (2019) to make it culturally relevant to Indonesian and Sundanese. We employ a multi-stage approach to adapt the data. First, we remove data that contains ambiguous options and offensive keywords. Next step involves assessing three elements: concept, name, and location, in order to sample the data that will be adapted. We take the question concepts from the existing CommonsenseQA data. As for the name and location extraction, we use Stanford’s Stanza NLP library Qi et al. (2020) and ConceptNet Speer et al. (2017) API. We use GPT-3.5 Turbo results ensembled from five different prompts to determine the relevance of the data based on the three elements. Data that are considered irrelevant in at least one of these three elements are selected to be adapted.
The next step is to transform the selected irrelevant data. First, we prompt GPT-4 Turbo [^3] Our initial experiment (Appendix A.2) reveals that GPT-4 Turbo significantly outperforms Merak-v4, an open Indonesian LLM, leading us to select GPT-4 Turbo for our work. to rephrase the sampled questions and options to align with Indonesian cultures. Subsequently, for data flagged with Western-centric names, a second rephrasing is conducted using the GPT-3.5 Turbo to replace all person names. The fully rephrased and adapted data are then translated from Indonesian to Sundanese using Google Translation API. Finally, we filter the data by examining the appearance of concepts in the questions and checking Sundanese-Indonesian back-translation with a similarity threshold of 90%.
@@ 4.2 Manual Data Generation
We employ annotators with different ethnicities from diverse regions covering Java and Bali island to ensure the variability of the knowledge in the human-generated dataset. The data collection process involves two main phases: (1) creating commonsense question-answer pairs and (2) answering commonsense questions. We perform quality control with additional human annotators to ensure the data quality.
@@
We first instruct the annotators to create commonsense question-answer pairs based on the given category and question concept. We also ask them to rely on their existing knowledge when making questions and avoid using internet search or LLM as much as possible. The categories and question concepts are manually curated to ensure wide domain coverage and to enrich the diversity of commonsense knowledge in the data. In total, we cover five categories: (1) culinary, (2) place, (3) culture, (4) history, and (5) activity, with 150 predefined question concepts for each category. Two annotators use the same 50 question concepts from each category to generate the questions, options, and answers. This process produces 1,500 QA pairs for each language in total.
@@
After passing the quality assurance, the 1,500 QA pairs are then redistributed among all annotators for further review, where they are tasked with answering 1,250 commonsense questions each (excluding the set of data they have made in the first phase). This phase aims to evaluate the consistency and variance in the commonsense knowledge among annotators from various cultural backgrounds. Annotators are instructed not only to provide answers to the questions but also to comment on any ambiguities in the questions or options, or if they have any uncertainties when answering the questions.
@@
We conduct Quality Control (QC) to ensure data correctness through manual human validation. QC annotators manually review data for errors and provide feedback, which is then corrected by the annotator concerned. Once corrected, QC annotators re-check the data to determine whether it can be considered complete or still requires revision. An evaluation meeting with all annotators is also held to convey a more comprehensive evaluation regarding all types of errors and other findings, ensuring annotators have a better understanding and do not repeat similar errors. After the second phase (answering questions), we also excluded some questions if more than three annotators marked them as ambiguous.
@@ 4.3 Automatic Data Generation
Dataset Version Indonesian Sundanese
Train Valid Test Train Valid Test
LLM_Adapt 1,506 191 158 1,506 191 158
Human_Gen - - 1,498 - - 1,499
LLM_Gen - - 1,063 - - 1,183
Total 1,506 191 2,719 1,506 191 2,840
4.4 Resulting Data
Our final CommonsenseQA dataset contains 4,416 QA pairs for Indonesian and 4,537 QA pairs for Sundanese. Table 1 provides detailed statistics for both languages. We also have three data variations: (1) LLM_Adapt: LLM-generated dataset constructed through data adaptation method (§4.1); (2) Human_Gen: Human-generated dataset, as explained in §4.2; (3) LLM_Gen: LLM-generated dataset given the same set of categories and concepts in Human_Gen4.3).
@@ 5.1.1 Concept Analysis
In the English (eng) to Indonesian (ind) concept adaptation, 95.57% are correctly adapted, including 16 out of 19 concepts needing major adaptation for cultural relevance, such as ‘snow’ to ‘hujan abu vulkanik’ (volcanic ashfall). Most adaptations are minor, primarily involving direct translations of eng to ind, as shown in Table 2. For Indonesian to Sundanese, the correct adaptation rate drops to 77.22%, reflecting weaker machine translation performance for Sundanese. The samples are shown in Table 3.
@@
Table 2: Samples of English (eng) to Indonesian (ind) concept adaptation.
Status Num (%) Concept Example
Orig (ind) Modified (sun)
Correct 122 (77.22%)
hujan abu vulkanik
(volcanic ashfall)
hujan lebu vulkanik
(volcanic ashfall)
menara (tower) munara (tower)
Wrong 36 (22.78%)
cicak
(house gecko)
kadal imah
(house lizard)
klinik gigi
(dental clinic)
klinik dental
(dental clinic)
@@
Despite the high accuracy of concept adaptation, as shown in Figure 2, we observe that the adapted concepts are skewed towards ‘komodo’. This indicates that the model is biased toward a specific entity within a category, which could arise from its limited knowledge or the absence of a direct equivalent for the original English concept in Indonesian, leading it to default to one standard concept. This finding highlights the need to manually develop concepts across various categories to improve the dataset’s quality in terms of concept diversity and coverage.
@@ 5.1.2 Question Analysis
To evaluate the quality of the generated questions, we apply a strict criterion: any errors, even minor, are marked as incorrect. As shown in Table 4, the Indonesian datasets show a high percentage of error-free questions, ranging between 68–75%. However, for Sundanese, the accuracy decreases significantly. In particular, the weak performance of the MT system in Sundanese is evident from its very low accuracy in LLM_Adapt. This suggests that synthetic data adaptation and generation is not useful for low-resource languages.
@@
To further analyze the quality of the questions, we delve deeper into the mistakes that are commonly made, which are listed in Table 5. We observe that most of the errors in LLM_Adapt come from translation errors. For LLM_Gen, even though the accuracy of question generation is lower than LLM_Adapt, 54.26% of the errors in LLM_Gen are actually minor typos or mechanics errors. In the case of Sundanese, most errors in LLM_Gen are due to sentence fluency, indicating that the model is still struggling to generate fluent sentences in Sundanese.

### v2

@@ Abstract
Can LLM Generate Culturally Relevant Commonsense QA Data?
Case Study in Indonesian and Sundanese
Rifki Afina Putri1, Faiz Ghifari Haznitrama1, Dea Adhista2, Alice Oh1 1KAIST, Republic of Korea 2Independent Researcher {rifkiaputri,haznitrama}@kaist.ac.kr, deadhista@gmail.com, alice.oh@kaist.edu
@@ 1 Introduction
The development of Large Language Models (LLMs) is significantly impacting NLP, leading to an increasing trend in the automated generation of datasets, particularly for question answering (QA) tasks. However, a major challenge arises with underrepresented languages like Indonesian and Sundanese due to the need for cultural context. For the generated data to be fully useful, it must not only be linguistically accurate, but it also needs to reflect the cultural nuances, historical references, and social norms. It is not yet clear whether current LLMs can create QA data that adequately includes the cultural nuances specific to certain languages.
Another common method for constructing non-English datasets is by utilizing a machine translation model. Although more scalable, this method cannot be straightforwardly applied due to the contextual irrelevancy of the data, primarily due to geographical differences (which could also influence cultural differences). For example, many English CommonsenseQA questions Talmor et al. (2019) include concepts such as snow or any winter sports, which are irrelevant in Indonesia due to its tropical climate with only two seasons. This dataset also often includes English-centric names and locations, primarily limited to the US, which are not considered “commonsense” for Indonesians.
@@
Therefore, in this study, we aim to investigate how well current LLMs can adapt and generate a commonsense QA dataset that is both linguistically accurate and culturally relevant to Indonesia. We focus on Indonesian, the lingua franca of Indonesia, and Sundanese, one of the regional languages in Indonesia with 34 million speakers, yet still considered low-resource Aji et al. (2022). To address dataset scarcity, especially in Sundanese, we also manually constructed datasets by involving annotators from various regions in Indonesia, ensuring a representation of diverse cultural perspectives. To sum up, our contributions are as follows:
[…]
  • We perform a comprehensive analysis to assess the effectiveness of LLMs in creating a culturally relevant commonsense QA dataset. We find that adaptation from English data is less effective, especially for Sundanese. However, GPT-4 Turbo can generate questions with some basic local knowledge in Indonesian and Sundanese when provided with human-created categories and concepts.
  • We evaluate several LLMs on our dataset and find that they perform better on LLM-generated data than on human-generated data, indicating that the former is less challenging, especially for proprietary models like GPT-4. Nevertheless, many open-source LLMs still struggle to answer LLM-generated questions, highlighting significant room for improvement in these datasets.
@@ 2 Related Work
Datasets for commonsense reasoning are crucial for developing models that understand and reason about real-world complexities. Datasets like COPA Roemmele et al. (2011), X-COPA Ponti et al. (2020), and The Winograd Schema Challenge Levesque et al. (2012) evaluate causal reasoning in real-world scenarios. Other datasets like ARC Clark et al. (2018), OpenBookQA Mihaylov et al. (2018), and Mcscript Ostermann et al. (2018) test commonsense reasoning via QA, but some questions require grade-school science knowledge. CommonsenseQA Talmor et al. (2019) presents the type of “purely” commonsense QA in the form of multiple-choice questions based on knowledge from ConceptNet Speer et al. (2017), built through crowdsourcing. However, the crowdworkers bring their cultural background as part of their common knowledge in building the data, resulting in many questions that are “commonsense” only in the Western culture.
@@
Various datasets are used to assess language models’ ability to understand cultural nuances specific to a language. FORK Palta and Rudinger (2023) explores culinary cultural biases and assumptions of US, Indian, and Chinese customs. IndoMMLU Koto et al. (2023) includes questions from Indonesian exams and covers regional cultural topics, such as the Minangkabau or Sundanese cultures. COPAL-ID Wibowo et al. (2023) is a COPA-style dataset written by native speakers, thus incorporating more Indonesian cultural nuances compared to X-COPA. COPAL-ID is composed of approximately 500 questions, with a primary emphasis on the cultural aspects of the Jakarta region. In contrast, our dataset is much bigger and covers a broader range of annotators from various regions of Java and Bali, in addition to Jakarta. We also include Sundanese, addressing the gap in this low-resource language.
@@ 3.1 Commonsense QA: Definition and Scope
In our dataset, we focus on building questions that probe common or cultural knowledge in daily life within Indonesian and Sundanese contexts. Each data point is a triple of concept, question, and options, with one correct answer. To enhance the relevance of the dataset to local contexts and reduce Western cultural bias, we not only utilized existing question concepts from the English CommonsenseQA Talmor et al. (2019) but also manually created novel question concepts. We cover five categories: (1) culinary, (2) place, (3) culture, (4) history, and (5) activity, with 150 unique question concepts for each category, aiming for broad domain and knowledge coverage.
@@ 3.2 Languages in Indonesia
Indonesia is one of the most culturally and linguistically diverse countries, with more than 700 languages spoken across the country Aji et al. (2022); Eberhard et al. (2021). Among the many languages, Indonesian is a unifying language used nationally. It utilizes the Latin script and was developed from literary "Classical Malay" Sneddon (2003), with regional variations. Over 80% of Standard Malay’s vocabulary is similar to Indonesian.
Apart from Indonesian, regional languages like Sundanese are spoken by people of the same ethnicity. It is primarily spoken in West Java and is the second-largest regional language in Indonesia, with 34 million speakers Eberhard et al. (2021). Regional languages, including Sundanese, have influenced the formation and development of the Indonesian language. Both languages share similarities, such as their grammatical structure, but also differ significantly in aspects like the number of vowels (i.e., Sundanese has 2 additional vowels: é, eu) and morphological features, including affixes.
We chose to study the national language and one regional language to illustrate the differences in the commonsense QA data generated via LLMs. It is also important to note that despite its large number of speakers, Sundanese has very limited data, and to date, there are no commonsense QA datasets available in Sundanese.
@@ 4 Data Generation Methods
To investigate whether LLMs can generate culturally relevant commonsense QA data in Indonesian and Sundanese, we construct a dataset using various methods with LLMs as data generators. In addition, we employ humans to generate data for comparison. As illustrated in Figure 1, we apply three dataset generation methods: (1) Automatic Data Adaptation (LLM_Adapt), where we leverage LLMs to automatically adapt English CommonsenseQA data to our target languages; (2) Manual Data Generation (Human_Gen), where we ask native-speaker human annotators to manually construct the dataset; and (3) Automatic Data Generation (LLM_Gen), where we use LLMs to generate data based on the list of categories and concepts used in Human_Gen. As shown in Table 1, our final dataset contains 8,953 QA pairs in total (4,416 for Indonesian and 4,537 for Sundanese).
Dataset Version Indonesian Sundanese
Train / Valid / Test Total Train / Valid / Test Total
LLM_Adapt 1,506 / 191 / 158 1,855 1,506 / 191 / 158 1,855
Human_Gen 0 / 0 / 1,498 1,498 0 / 0 / 1,499 1,499
LLM_Gen 0 / 0 / 1,063 1,063 0 / 0 / 1,183 1,183
1,506 / 191 / 2,719 4,416 1,506 / 191 / 2,840 4,537
While all prompts given to the LLM are in English, for the LLM_Adapt data, we specifically instruct the model to produce responses directly in Indonesian. Subsequently, we translate the resulting Indonesian data into Sundanese. In contrast, for the LLM_Gen data, we instruct the model to give the output directly in Indonesian and also Sundanese, to closely replicate the data generation process used for Human_Gen. The generation methods for each dataset are detailed below.
@@ 4.1 Automatic Data Adaptation
We build the first LLM-generated data by adapting the CommonsenseQA dataset Talmor et al. (2019) to make it culturally relevant to Indonesian and Sundanese. First, we remove data that contains ambiguous options and offensive keywords. Next step involves assessing three elements: concept, name, and location, in order to sample the data that will be adapted. We take the question concepts from the existing CommonsenseQA data. As for the name and location extraction, we use Stanford’s Stanza NLP library Qi et al. (2020) and ConceptNet Speer et al. (2017) API. We use GPT-3.5 Turbo results ensembled from five different prompts to determine the relevance of the data based on the three elements. Data that are considered irrelevant in at least one of the three elements are selected to be adapted. The examples are shown in Table 11.
The next step is to transform the selected irrelevant data. First, we prompt GPT-4 Turbo [^2] Our initial experiment (Appendix A.2) reveals that GPT-4 Turbo significantly outperforms Merak-v4, an open Indonesian LLM, leading us to select GPT-4 Turbo for our work. to rephrase the sampled questions and options to align with Indonesian cultures. Subsequently, for data flagged with Western-centric names, a second rephrasing is conducted using the GPT-3.5 Turbo to replace all person names. The fully rephrased and adapted data are then translated from Indonesian to Sundanese using Google Translation API. Finally, to ensure the dataset quality, we filter the data by examining the appearance of concepts in the questions and checking Sundanese-Indonesian back-translation with a similarity threshold of 90%.
@@ 4.2 Manual Data Generation
To build the human-generated dataset, we recruited 12 experienced annotators from diverse regions across Java and Bali, as detailed in Table 10. All are native speakers of Indonesian (and Sundanese), with a minimum of 15 years residing in their particular target language regions. Among them, 5 annotators have a degree in linguistics. These criteria ensure they possess both language proficiency and cultural familiarity. As for the data collection process, we have two main phases: (1) creating commonsense question-answer pairs and (2) answering commonsense questions. We also perform quality control to ensure the data quality. [^3] More details on annotators’ demographic information are presented in Appendix B. The complete annotation guideline is outlined in Appendix C.
@@
We first instruct the annotators to create commonsense question-answer pairs based on the given category and question concept that we had newly created (§3.1). We also ask them to rely on their existing knowledge when making questions and avoid using internet search or LLM as much as possible. Two annotators use the same 50 question concepts from each category to generate the questions, options, and answers. This process produces 1,500 QA pairs for each language in total.
@@
After passing the quality assurance, the 1,500 QA pairs are then redistributed among all annotators for further review, where they are tasked with answering 1,250 commonsense questions each (excluding the set of data they have made in the first phase). Therefore, in this phase, each question is answered by five different annotators to capture the consistency and variance in the commonsense knowledge among annotators from various cultural backgrounds. Annotators are instructed not only to provide answers to the questions but also to comment on any ambiguities in the questions or options, or if they have any uncertainties when answering the questions.
@@
We conduct Quality Control (QC) through manual validation to ensure data accuracy and maintain its “commonsense” nature. QC annotators manually review data for errors and provide feedback, which is then corrected by the annotator concerned. Once corrected, QC annotators re-check the data to determine whether it can be considered complete or still requires revision. An evaluation meeting with all annotators is also held to convey a more comprehensive evaluation regarding all types of errors and other findings, ensuring annotators do not repeat similar errors and have a better understanding of the notion of commonsense. After the second phase (answering questions), we also excluded some questions if more than three annotators marked them as ambiguous.
@@ 5.1.1 Concept Analysis
In the English to Indonesian concept adaptation, 95.57% are correct, including 16 out of 19 concepts needing major adaptation for cultural relevance, such as ‘snow’ to ‘hujan abu vulkanik’ (volcanic ashfall), showing that the LLM has some basic knowledge to adapt English concepts, although most adaptations primarily involving exact translations without major alteration of the concept (Table 2). For Indonesian to Sundanese, the correct adaptation drops to 77.22%, reflecting weak machine translation performance in Sundanese (Table 8).
@@
Table 2: eng to ind concept adaptation result.
@@
Despite the high accuracy of concept adaptation, as shown in Figure 2, the adapted concepts are skewed towards ‘komodo,’ indicating a bias towards a specific entity within a category. This could be due to the model’s limited knowledge or the absence of a direct equivalent for the original English concept in Indonesian, leading it to default to one standard concept. This finding highlights the limitations of adaptation and translation methods, and to improve dataset diversity and coverage, we need to manually develop concepts across various categories.
@@ 5.1.2 Question Analysis
To evaluate the quality of the generated questions, we apply a strict criterion: any errors, even minor, are marked as incorrect. As shown in Table 3, the Indonesian datasets show a high percentage of error-free questions, ranging between 68–75%. However, for Sundanese, the accuracy decreases significantly. In particular, the weak performance of the MT system in Sundanese is evident from its very low accuracy in LLM_Adapt, suggesting synthetic data adaptation/translation from English is not an optimal method for low-resource languages. In contrast, the LLM_Gen dataset, which involves generating synthetic data directly in the target language, shows a significantly higher number of correct questions, particularly for Sundanese. This indicates that direct generation in the target language is a more promising approach than adaptation or translation.
@@
We manually reviewed the questions to identify common errors, detailed in Table 5.1.3. We observe that most of the errors in LLM_Adapt come from translation errors. For Indonesian LLM_Gen, despite lower question generation accuracy, most of the errors are actually minor typos. In Sundanese LLM_Gen, however, the predominant issues relate to sentence fluency. This indicates that the model is still struggling to generate fluent sentences in Sundanese, where it occasionally uses phrases or words in Indonesian instead of Sundanese. While Indonesian and Sundanese share many linguistic features, they differ significantly in morphological features (§3.2). This characteristic, along with the smaller amount of Sundanese data used for training LLMs, results in the models being less exposed to Sundanese words, suffixes, or prefixes compared to Indonesian. This leads to the occasional appearance of Indonesian words in Sundanese data, even though the sentences may be grammatically correct.